OpenStudy (anonymous):

Which is an equation of the line with slope -3/4 passing through the point (-,2,-10)? A. y = -3/4x -23 B. y = -3/4x = 19/2 C. y = -3/4x - 10 D. 3x - 4y = 23 Please give me step by step, thank you!

OpenStudy (fibonaccichick666):

ok, so, fastest way is to just plug your x value in and see which gives you the required y

OpenStudy (fibonaccichick666):

can you try that please?

OpenStudy (anonymous):

Slope intercept form would be y=mx+b (slope is m and b is y intercept) slope: -3/4 y intercept: -10 Put it together! y=-3/4x-10 so the answer would be C Hope it helps!

OpenStudy (anonymous):

@Breezy6 please read: http://openstudy.com/code-of-conduct thanks

OpenStudy (anonymous):

Actually I didn't just GIVE the answer, I explained it as the rules state. "Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem."

OpenStudy (fibonaccichick666):

I agree with Data. That is the same as giving an answer.

OpenStudy (anonymous):

you should also encourage them to join solving their own problems because that's the only way they will learn.

OpenStudy (anonymous):

So, the b is -10? Because it's the Y in the coordinates? :)

OpenStudy (anonymous):

Yes, the y intercept is the b. y=mx+b m= slope b= y intercept

OpenStudy (fibonaccichick666):

uhm, no, so in order for the -10 to be b, it would have to be the y-intercept. Since you do not have that, you can either come up with an equation of your own and compare or plug and chug

OpenStudy (anonymous):

Slope: 1/2 (8, -6) Y = 1/2x + -6 Am I correct?

OpenStudy (anonymous):

Y = 1/2x - 6**

OpenStudy (fibonaccichick666):

what? that's not even related to this problem

OpenStudy (anonymous):

look at the third post..?

OpenStudy (anonymous):

I know, its the next one, Im using it to see If I understood what he told me

OpenStudy (anonymous):

Breezy is a she btw and what is your next problem?

OpenStudy (fibonaccichick666):

No, so in order to use the y coordinate as b, your x coordinate must =0

OpenStudy (anonymous):

Which is an equation of the line with slope 1/2 passing through the point (8,-6)?

OpenStudy (fibonaccichick666):

so currently, you just have a point and a slope

OpenStudy (anonymous):

So, why did she use -10 as the B in the first one? x was -2

OpenStudy (fibonaccichick666):

the best way to get the eq. is to use Point-slope form, do you know it?

OpenStudy (anonymous):

Y = mx + or - b?

OpenStudy (fibonaccichick666):

because it is wrong, so ignore that

OpenStudy (fibonaccichick666):

please, and I will show you the correct way

OpenStudy (anonymous):

Is y = mx + or - b, Point-slope form?

OpenStudy (anonymous):

You can use a given point and a slope to find a line

OpenStudy (anonymous):

y=mx+b you are given the slope (m) to find the value of "b" you should use the given point and substitute it in the equation y=mx+b and solve for b.

OpenStudy (fibonaccichick666):

breezy please stop you are incorrect

OpenStudy (anonymous):

Ok, data let me do this

OpenStudy (fibonaccichick666):

@kmcllb you want to use point slope form for this problem

OpenStudy (anonymous):

That's what I literally said twice. "Yes, the y intercept is the b. y=mx+b m= slope b= y intercept"

OpenStudy (anonymous):

Is the answer y = 1/2x -2?

OpenStudy (fibonaccichick666):

sounds more likely, but can I see your work please?

OpenStudy (anonymous):

-6 = 1/2(8) +b -6 = 4 + b -4 -4 -10 = b y = 1/2x - 10

OpenStudy (fibonaccichick666):

very good. Yes, that is correct

OpenStudy (anonymous):

Thank you, :)

OpenStudy (fibonaccichick666):

now you can also use, \(y-y_0=m(x-x_0)\) when \((x_0,y_0)\) is any point on the line

OpenStudy (anonymous):

I have another question

OpenStudy (fibonaccichick666):

that is called point slope form, and it tends to be more useful. Sure just tag me

OpenStudy (anonymous):

Which is an equation of the line in the figure (I have no figure) I can tell you the points (0,4) (0,-3)

OpenStudy (anonymous):

@fib

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

A. 3x + 4y = 0 B. 4x + 3y = o C. 4x + 3y = 12 D. 3x +4y = 12

OpenStudy (fibonaccichick666):

ok, so, first, you need to find the slope. If you have two points, do you know how to find slope?

OpenStudy (fibonaccichick666):

Although, your question doesn't make sense, did you copy it correctly?

OpenStudy (anonymous):

Yes, y^2 - y^1 x^2 - x^1

OpenStudy (anonymous):

Can I insert a picture taken from an iPhone on here to show you my Booklet?

OpenStudy (fibonaccichick666):

That is slope form, I assume you mean a fraction there, but (0,4) (0,-3), are those the points? That's my issue, can you just check those?

OpenStudy (fibonaccichick666):

slope formula*

OpenStudy (anonymous):

There are no point's, you have to choose your own, by looking at the line (It's a negative slope if that helps)

OpenStudy (fibonaccichick666):

ohhhh ok, so you cannot have two points with an x coordinate of zero and have a line.

OpenStudy (fibonaccichick666):

so, do you know how to find the y value when x=0?

OpenStudy (anonymous):

(0,4) (3,0) Did I say it wrong?

OpenStudy (fibonaccichick666):

yea, ok that sounds better

OpenStudy (fibonaccichick666):

now, make sure those are of this form (x,y)

OpenStudy (fibonaccichick666):

Then, calculate your slope

OpenStudy (anonymous):

(0,4) (3,0) 0 - 4 = -4 3 - 0 = 3 4 = -4/3(0) + b 4 = 0 + b -0 -0 4 = b y = -4/3x + b?

OpenStudy (anonymous):

NO!

OpenStudy (fibonaccichick666):

ok, fix it

OpenStudy (anonymous):

4x + 3y = 0

OpenStudy (anonymous):

In standard form

OpenStudy (fibonaccichick666):

no, better with last try. What is slope formula again?

OpenStudy (anonymous):

It need's to be in standard form

OpenStudy (fibonaccichick666):

you're getting ahead of yourself

OpenStudy (anonymous):

Y = Mx + n My answer was: y = -4/3 + 4 It's not an asnwer choice, so i need to make it standard form

OpenStudy (anonymous):

y = mx + b

OpenStudy (fibonaccichick666):

That is standard form, but your slope is wrong

OpenStudy (fibonaccichick666):

hence, go back and tell me the slope formula again

OpenStudy (anonymous):

0 - 4 -4 3 - 0 3

OpenStudy (anonymous):

y2 - y1 x2 - x1

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

4-0 3-0 4/3

OpenStudy (anonymous):

No

OpenStudy (anonymous):

(0, 4) (3, 0) y2 -y1 x2 - x1 0-4 -4 3-0 3 Am I wrong?

OpenStudy (fibonaccichick666):

uhm well which is your answer?

OpenStudy (anonymous):

Well, -4 Divided by 3 is -1.333 OR -4/3

OpenStudy (fibonaccichick666):

that is correct, now, what is your y intercept?

OpenStudy (anonymous):

4

OpenStudy (fibonaccichick666):

good, so now, what is your slope intercept form?

OpenStudy (fibonaccichick666):

(btw, slope intercept is usually called standard form, at least in my experience, not that blah they are making you do)

OpenStudy (anonymous):

Y = -4/3 + 4 Standard form: Ax + By = c 4x + 3y = 4, which is not one of my answer choices

OpenStudy (fibonaccichick666):

no, you haven't done it yet

OpenStudy (fibonaccichick666):

so first, fx your slope intercept. It's missing something

OpenStudy (anonymous):

Slope intercept?

OpenStudy (anonymous):

My equation?

OpenStudy (fibonaccichick666):

the formula you gave is wrong, it's missing stuff

OpenStudy (anonymous):

y2 - y1 x2 -x1

OpenStudy (anonymous):

y = mx + or - B

OpenStudy (fibonaccichick666):

No, wait, I guess I wasn't clear. BTW, it's just y=mx+b there is no minus. b can be negative though

OpenStudy (anonymous):

I know, thats why I say + or - Ex. Y = 1x - 1 Ex. Y - 1x + 1

OpenStudy (fibonaccichick666):

Y = -4/3 + 4<-----I'm missing something important Standard form: Ax + By = c 4x + 3y = 4, which is not one of my answer choices

OpenStudy (anonymous):

Y = -4/3x + 4

OpenStudy (fibonaccichick666):

But it is actually incorrect to say plus or minus in the formula. It implies two answers

OpenStudy (fibonaccichick666):

ok, now, how can you get the x to the other side?

OpenStudy (anonymous):

By subtracting it

OpenStudy (anonymous):

Nononon, Dividing by it

OpenStudy (fibonaccichick666):

ok, subtracting what exactly?